### v2

# COMPAS: A Distributed Multi-Party SWAP Test for Parallel Quantum Algorithms

Abstract.
The limited number of qubits per chip remains a critical bottleneck in quantum computing, motivating the use of distributed architectures that interconnect multiple quantum processing units (QPUs). However, executing quantum algorithms across distributed systems requires careful co-design of algorithmic primitives and hardware architectures to manage circuit depth and entanglement overhead. We identify multivariate trace estimation as a key subroutine that is naturally suited for distribution, and broadly useful in tasks such as estimating Rényi entropies, virtual cooling and distillation, and certain applications of quantum signal processing. In this work, we introduce COMPAS, an architecture that realizes multivariate trace estimation across a multi-party network of interconnected modular and distributed QPUs by leveraging pre-shared entangled Bell pairs as resources. COMPAS adds only a constant depth overhead and consumes Bell pairs at a rate linear in circuit width, making it suitable for near-term hardware. Unlike other schemes, which must choose between asymptotic optimality in circuit depth or GHZ width, COMPAS achieves both at once. Additionally, we analyze network-level errors and simulate the effects of circuit-level noise on the architecture. [^1] The code is available at github.com/kunliu7/Distributed-Q-Algo.
Quantum computing; Distributed architectures; Entanglement distillation
isbn: 978-1-4503-XXXX-X/2018/06journalyear: 2026copyright: ccconference: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 22–26, 2026; Pittsburgh, PA, USAbooktitle: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’26), March 22–26, 2026, Pittsburgh, PA, USAdoi: 10.1145/3779212.3790143isbn: 979-8-4007-2359-9/2026/03ccs: Hardware Quantum computationccs: Computer systems organization Distributed architectures
1. Introduction
Scaling physical quantum systems while maintaining high qubit connectivity and gate fidelity presents significant challenges due to constraints in control complexity, wiring density, power bottlenecks, and maintenance overhead. Parallel and distributed quantum computing (margolus1990parallel; caleffi_2024_distributed; beals_efficient_2013) have emerged as two promising paradigms to address these limitations by leveraging modular designs that interconnect smaller quantum processors via quantum communication channels. However, harnessing such modular architectures demands careful co-design of quantum algorithms and the underlying hardware to overcome the fundamental tension between locality, latency, and fidelity (kim_fault-tolerant_2024; mor-ruiz_influence_2024; NoisyDistributedQC). One fundamental challenge is how to perform efficient remote entangling operations across multiple parties.
In distributed quantum computing, efficient execution of algorithms relies on high-rate and low-latency communication between modules (parekh_quantum_2021; Cacciapuoti2020quantum). The basic resource for such a modular system is a Bell pair shared between physical qubits in different nodes, enabling quantum-state teleportation and remote gate operations. Developing an interface with sufficient rate and fidelity to support high-fidelity computation remains a significant challenge. There has been tremendous experimental progress developing interconnections among quantum systems, such as using cryogenic microwave links for superconducting qubits or optical photons for atomic qubits (ang2024arquin; main2025distributed). Yet, state-of-the-art results for remote entanglement between neutral atom or trapped ion qubits only achieves rates of a few hundred Hz with entangling fidelities in the high 90% (Stephenson_2020; jing2019entanglement; Ritter_2012; Hucul_2014; Pompili_2021). Entanglement distillation protocols have also been developed to further boost the fidelity of entanglement while maintaining a high rate and low latency (rozpkedek2018optimizing). Despite this progress, architectures that can efficiently connect modular and distributed quantum hardware with generic high-level quantum algorithms are much underdeveloped.
This paper introduces COMPAS (Compiling a Multi-Party SWAP Test for Parallel Algorithms on Distributed Quantum Systems) as a co-designed hardware/software framework for executing multivariate trace estimation, a core quantum primitive, across distributed quantum systems. At the heart of COMPAS is the key insight that the program structure of these algorithms permits a structured decomposition of computation and communication: local operations can be kept shallow, while remote gates can be efficiently handled using Bell pair assisted teleportation.
More specifically, COMPAS performs a particular multivariate trace estimation known as the multi-party SWAP test, which underlies a wide range of parallel algorithms and protocols in quantum computing. Unlike general quantum circuits, the multi-party SWAP test has a highly structured design, which presents an opportunity for targeted compiler optimizations that minimize communication costs while preserving algorithmic running time and fidelity. Distributed quantum circuits must simultaneously account for multiple conflicting constraints, including spatial limitations (qubit count), temporal efficiency (circuit depth), and inter-node communication (entanglement distribution and non-local operations). Logical circuits designed for monolithic architectures often do not translate efficiently to distributed systems, necessitating specialized compiler techniques to minimize long-range operations and reduce interconnect overhead. Without deep optimization, the resource overhead associated with parallel quantum algorithms can quickly become prohibitive, negating the theoretical advantages of parallel execution. A tailored compilation approach must consider qubit routing, gate scheduling, and efficient use of entanglement to ensure that logical operations can be executed with minimal additional cost while maintaining high fidelity. COMPAS provides an architecture design that naturally separates the individual pieces of data involved in the multi-party SWAP test into distinct modules. This reduces both the amount of network communication required, because state preparation can be performed within individual modules, and improves synchronization and parallelism, because the preparation of these inputs can happen in parallel across different modules.
In presenting COMPAS, we develop techniques to adapt the Quek et al. implementation of the multi-party SWAP test (quek_2024) into a distributed quantum algorithm and architecture, while maintaining constant circuit depth independent of the number of modules. The constant-depth property is crucial as it ensures that the runtime is not slowed down by communication between nodes, thereby enabling true parallelism at scale. Our approach optimizes both qubit usage and inter-node communication, while also remaining robust to noisy communication and gate errors, which we confirm via circuit-level simulation. We further highlight applications of COMPAS to a wide range of computation tasks, including Rényi entropy calculation (Yirka_2021; Islam_2015), entanglement spectroscopy (Johri_2017), virtual cooling and distillation (Huggins_2021; Cotler_2019), and parallel quantum signal processing (martyn_2024_parallel). In aggregate, by carefully balancing circuit design, entanglement distribution, and hardware constraints, we provide a pathway toward scalable, high-performance quantum computing on distributed architectures.
1.1. Results and Paper Outline
We begin with an overview of the necessary background in Sec. 2, reviewing methods in distributed quantum computing and the multi-party SWAP test. We then describe our proposed architecture in Sec. LABEL:sec:approach. A central contribution of this section is a distributed implementation of the controlled-SWAP (CSWAP) gate, which is a key operation used in recent multivariate trace estimation algorithms (quek_2024). In Secs. LABEL:sec:telegate_design and LABEL:sec:teledata_design, we present two such approaches—one based on the teledata primitive, and the other on the telegate primitive, the two core tools of distributed quantum computing. In Sec. LABEL:sec:resources, we analyze the resource costs of the proposed architectures, including circuit depth, and entanglement consumption. We simulate and analyze the impact of gate-level noise in Sec. LABEL:sec:error_analysis, and finally highlight applications of the proposed architecture to various quantum algorithms in Sec. LABEL:sec:applications.
2. Background
2.1. High-Level Overview
[…]
The multi-party SWAP test is a standard tool for comparing many quantum states. It operates by extracting the trace of a product of states, a quantity in the interval $[0,1]$ that provides information about the states’ overlaps, spectra, and entanglement structure. However, implementing the multi-party SWAP test on near-term hardware is expensive due to hardware limitations. In spite of this, distributed quantum computing offers a potential remedy by distributing computational tasks across a network of interconnected quantum processing units (QPUs). This reduces the computational load on each device, but can incur additional communication overhead.
In this work, we present a distributed implementation of the multi-party SWAP test that simultaneously achieves optimal latency and memory usage. To our knowledge, this is the first approach to do so. Our construction, and distributed quantum protocols in general, use pairs of correlated (entangled) states known as Bell pairs, which are distributed between parties in advance. These pre-shared resources enable spatially separated QPUs to interact with each other without directly transmitting quantum states during runtime. We leverage this capability to distribute the SWAP test over many parties while consuming Bell pairs as a resource.
2.2. Distributed Quantum Computing
In this work, we distinguish between parallel and distributed quantum algorithms. We use parallel quantum algorithms to refer to algorithms that reduce circuit depth at the expense of increased qubit count. Such parallelism may occur on a single quantum processor or across multiple processors.
In contrast, distributed quantum algorithms are designed to partition a computational task across multiple spatially separated QPUs, typically requiring classical or quantum communication between them. While having the potential to solve problems that exceed the resources of any single QPU (Cuomo_2020), distributed quantum computing faces obstacles not present in classical distributed computing. Notably, the no-cloning theorem forbids the copying of quantum states between processors. Instead, parties commonly interact through the teledata and telegate primitives (Fig. LABEL:fig:TeledataTelegate), which we collectively refer to as teleoperations. Each teleoperation leverages pre-shared Bell pairs and classical communication as resources. A teledata operation refers to the teleportation of a quantum state between parties, whereas telegate operations allow one party to act as the control for a unitary acting on another.
[…]